$SPAD/src/input f01rcf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)cl all
showArrayValues true
showScalarValues true
m:=5
n:=3
lda:=5
a:Matrix Complex SF:=
 [[complex(0,0.5) ,complex(-0.5,1.5),complex(-1.0,1.0)],_
  [complex(0.4,0.3),complex(0.9,1.3),complex(0.2,1.4) ],_
  [complex(0.4,0), complex(-0.4,0.4),complex(1.8,0)],_
  [complex(0.3,-0.4),complex(0.1,0.7),complex(0,0)],_
  [complex(0,-0.3),complex(0.3,0.3),complex(0,2.4) ]]
result:=f01rcf(m,n,lda,a,-1)

———-

References^{[1] nothing}

4